# Comparative Studies of Quantum Annealing, Digital Annealing, and Classical Solvers for Reaction Network Pathway Analysis and mRNA Codon Selection

### v1

@@ 5 Future work
The following work would be useful to evaluate:
  • Use cases with different kinds of QUBOs, such as problems with quadratic constraints, would be useful to evaluate. Problems with superlinear bloating due to binarization, or HOBO to QUBO mappings would also be interesting benchmarks but potentially more difficult for QUBO solvers.
  • Benchmark the D-Wave Leap NL HQA solver and CP-SAT for mRNA codon selection with larger problem sizes and with larger runtimes than were used in this work’s benchmarking of the Extra-large Proteins.
@@ 6 Conclusions
This work developed a benchmarking framework for evaluating the utility of quantum annealing, digital annealing, and classical solvers for combinatorial optimization problems formulated as QUBOs with constraints. The framework analyzes problem mapping metrics, quantitative connectivity measures, and penalty structures to systematically assess when quantum and quantum-inspired methods may offer advantages over classical approaches. Through benchmarking two industrially relevant use cases in chemistry and life sciences, reaction network pathway analysis and mRNA codon selection, we observed that problem structure influences performance of the different kinds of solvers. For reaction network pathway analysis problems, which exhibit linear/logarithmic bloating in QUBO formulation, low density, and linear constraints, classical solvers (Gurobi, HiGHS, SCIP, CP-SAT) outperformed digital annealing in both accuracy and efficiency. For mRNA codon selection, the problems had one-to-one mapping to QUBO, higher density and interconnectivity, one-hot constraints, however they also had high rank-1 dominance. CP-SAT solved the Standard and Large Protein Dataset problems (under 1500 amino acids) the most efficiently; the CP-SAT, Fujitsu DA, D-Wave Leap NL HQA, and HiGHS solvers achieved optimal solutions with linear computational scaling compared to exponential scaling with the D-Wave Leap CQM HQA and SCIP solvers. The D-Wave Leap CQM HQA which natively supports constraints found near-optimal solutions for all problems, and unconstrained D-Wave HQA solvers were unable to identify near-optimal solutions in reasonable amounts of time. In the Extra-large Protein Dataset (11000 to 14000 amino acids), the D-Wave NL HQA solver performed comparably to CP-SAT, outperforming it in minimum cost in 2 out of the 4 problems, out of which in one CP-SAT did not identify a feasible solution within the time limit. These findings highlight the importance of problem-specific evaluation and provide practical guidance for selecting appropriate optimization methods based on QUBO structure and constraint types.
@@ 7 Acknowledgements
The authors would like to thank Fujitsu for providing access to the digital annealer during a proof of concept project and the Fujitsu Accelerator Program, as well as Matthieu Parizy for his support during these programs. We would also like to thank QuantumBasel for providing access to D-Wave quantum annealers, and appreciate Julien Baglio for his advice in benchmarking these solvers. Additionally, we would like to thank Helmut Katzgraber for his advice in constraint programming and presenting benchmark results.

### v2

@@ 5 Future work
It would be useful to evaluate use cases with different kinds of QUBOs, such as problems with quadratic constraints and dense problems that do not have high rank-1 dominance. Problems with superlinear bloating due to binarization, or HOBO to QUBO mappings would also be interesting benchmarks but potentially more difficult for QUBO solvers.
@@ 6 Conclusions
This work developed a benchmarking framework for evaluating the utility of quantum annealing, digital annealing, and classical solvers for combinatorial optimization problems formulated as QUBOs with constraints. The framework analyzes problem mapping metrics, quantitative connectivity measures, and penalty structures to systematically assess when quantum and quantum-inspired methods may offer advantages over classical approaches. Through benchmarking two industrially relevant use cases in chemistry and life sciences, namely reaction network pathway analysis and mRNA codon selection, we observed that problem structure influences performance of the different kinds of solvers. For reaction network pathway analysis problems, which exhibit linear/logarithmic bloating in QUBO formulation, low density, and linear constraints, classical solvers (Gurobi, HiGHS, SCIP, CP-SAT) outperformed digital annealing in both accuracy and efficiency. For mRNA codon selection, the problems had one-to-one mapping to QUBO, higher density and interconnectivity, one-hot constraints, however they also had high rank-1 dominance. Gurobi solved all problems (up to 14000 amino acids) the most efficiently; the Gurobi, CP-SAT, Fujitsu DA, D-Wave Leap NL HQA, and HiGHS solvers achieved optimal solutions with linear computational scaling compared to exponential scaling with the D-Wave Leap CQM HQA and SCIP solvers. The D-Wave Leap CQM HQA which natively supports constraints found near-optimal solutions for all problems, and unconstrained D-Wave HQA solvers were unable to identify near-optimal solutions in reasonable amounts of time. In the Extra-large Protein Dataset (11000 to 14000 amino acids), Gurobi exhibited the best performance, and the D-Wave NL HQA solver performed comparably to CP-SAT. These findings highlight the importance of problem-specific evaluation and provide practical guidance for selecting appropriate optimization methods based on QUBO structure and constraint types.
@@ 7 Acknowledgements
The authors would like to thank Fujitsu for providing access to the digital annealer during a proof of concept project and the Fujitsu Accelerator Program, as well as Matthieu Parizy for his support during these programs. We would also like to thank QuantumBasel for providing access to D-Wave quantum annealers, and appreciate Julien Baglio for his advice in benchmarking these solvers. Additionally, we would like to thank Helmut Katzgraber (55 North) for his advice in constraint programming and presenting benchmark results.